a piece of string of length 85 centimeter is divided into three pieces in the ratio 2:3:5 calculate the lengthof the length of the shorteset and longest .
An alloy consist od steel silver and copper in the ratio of 6:5:9 . If the smallest masss is 160 grams , calculate the mass of the copper i the alloy

Found 3 solutions by rfer, addingup, ikleyn:
Answer by rfer(16322)

2x+3x+5x=85
10x=85
x=85/10
x=8.5
2x=17 shortest
5x=42.5 longest
---------------
6x+5x+9x=160
20x=160
x=160/20
x=8
9x=72 copper

Answer by addingup(3677)

The length is 85 and the ratio is 2:3:5. Now, 2+3+5 = 10, which means that in every group of 10 there are 2 of one, 3 of another, and 5 of yet another.
How many groups of 10 go into 85? Well, 85/10 = 8.5. Aha! We've got our answer:
(2 x 8.5) + (3 x 8.5) + (5 x 8.5) = 85 let's see if this is true:
17 + 25.5 + 42.5 = 85 Correct
-----------------------------------------
An alloy consist of steel silver and copper in the ratio of 6:5:9 . the smallest mass (silver) is 160 grams. So 5 = 160
IF:
5 = 160
then:
(160*6)/5 = 192
(160*9)/5 = 2886 x 160)/5 = 288
mass of silver: 160
mass of steel: 192
mass of copper: 288
-------------------------
The mass of copper in the alloy is 289
-----------------------------------------
:

Answer by ikleyn(52847)

.
The correct answer is

    The mass of copper in the alloy is 288 grams.


Any other answer is   I N C O R R E C T.

The solution by @rfer is   I R R E L E V A N T.